OpenStudy (anonymous):

Find the monthly payment needed to pay off a loan of $3900 amortized at 3% compounded monthly for 6 years.

OpenStudy (amistre64):

Pert

OpenStudy (amistre64):

3900 e^(.03)(6)

OpenStudy (anonymous):

The answer doesn't make sense though..

OpenStudy (amistre64):

3900 * (e^(.03 * 6)) = 4,669.14772 Divide that by 72 months to get the payments 64.85

OpenStudy (anonymous):

It says it's wrong :/

OpenStudy (amistre64):

64.84? let me look up the definition for amoritized...

OpenStudy (anonymous):

Yeah, I have this online homework called "Webwork" and it says that's incorrect.. i've already tried

OpenStudy (amistre64):

compunded monthly.....I did continuously becasue I cant read...im going illiterate

OpenStudy (anonymous):

It's okay! lol

OpenStudy (amistre64):

P(1+(.03/12))^(12*6) should be the total amount and then divide tha tby 72 to get the monthlies...

OpenStudy (amistre64):

3 900 * ((1 + (.03 / 12))^(12 * 6)) = 4,668.09902 divide by 72: 64.83 trythat :)

OpenStudy (anonymous):

Nope :/

OpenStudy (amistre64):

....... new tax laws maybe? :)

OpenStudy (anonymous):

haha Maybe.. it's getting tedious though

OpenStudy (amistre64):

P(1+r/n)^(n*t) is the formula. that gives you the total amount of the loan over t years. right?

OpenStudy (amistre64):

\[3900(1+\frac{.03}{12})^{12*6}\]

OpenStudy (anonymous):

Tried that already

OpenStudy (amistre64):

if it aint that....I dont know what id be..... amortized just means. 59.26 is what an amortization calculator online gives me////

OpenStudy (anonymous):

Thank you though (:

OpenStudy (anonymous):

wait could you please copy and paste that url.. the "amortization calculator online"

OpenStudy (anonymous):

Because I have one more just like this to do that webwork kept saying was wrong

OpenStudy (amistre64):

http://www.amortization-calc.com/

OpenStudy (anonymous):

what in the world haha

OpenStudy (amistre64):

gotta put in the right numbers :)

OpenStudy (anonymous):

ughhh lol thank you. this is difficult

OpenStudy (anonymous):

:)

OpenStudy (dumbcow):

this is more of a finance question a finance calculator like a BA II would help you a lot the math behind it deals with present value equations the value of your payment changes based on how much goes toward principal and how much towards interest since with every payment the balance decreases the interest amount will also decrease over time anyway the equations is this i = .03/12 = .0025 v = 1/(1+i) = .997506 3900 = P(1+v+v^2 +...v^71) sum of (1+v+...) = (1-v^72)/(1-v) = 65.98 P=3900/65.98 = 59.11 this is an approximate answer and prob more info than you need i put this into my BA II and got 59.25